# A Survey on Wi-Fi Sensing Generalizability: Taxonomy, Techniques, Datasets, and Future Research Prospects

### v2

@@ Abstract
Wi-Fi sensing has emerged as a powerful non-intrusive technology for recognizing human activities, monitoring vital signs, and enabling context-aware applications using commercial wireless devices. However, the performance of Wi-Fi sensing often degrades when applied to new users, devices, or environments due to significant domain shifts. To address this challenge, researchers have proposed a wide range of generalization techniques aimed at enhancing the robustness and adaptability of Wi-Fi sensing systems. In this survey, we provide a comprehensive and structured review of over 200 papers published since 2015, categorizing them according to the Wi-Fi sensing pipeline: experimental setup, signal preprocessing, feature learning, and model deployment. We analyze key techniques, including signal preprocessing, domain adaptation, meta-learning, metric learning, data augmentation, cross-modal alignment, federated learning, and continual learning. Furthermore, we summarize publicly available datasets across various tasks—such as activity recognition, user identification, indoor localization, and pose estimation—and provide insights into their domain diversity. We also discuss emerging trends and future directions, including large-scale pretraining, integration with multimodal foundation models, and continual deployment. To foster community collaboration, we introduce the Sensing Dataset Platform (SDP) for sharing datasets and models. This survey aims to serve as a valuable reference and practical guide for researchers and practitioners dedicated to improving the generalizability of Wi-Fi sensing systems.
@@
Wi-Fi sensing, Human sensing, Action recognition, Indoor localization, Domain adaptation, Domain generalization, Few-shot learning, Generative adversarial network, Transfer learning, Meta-learning
@@ I Introduction
While Wi-Fi sensing has shown great promise across a variety of applications, its effectiveness fundamentally relies on how wireless signals interact with the surrounding environment and the human body. Fig. 1 illustrates how human presence and motion influence the propagation of Wi-Fi signals. As shown in the figure, Wi-Fi devices are typically equipped with multiple transmit and receive antennas. Transmit antennas emit radio signals that propagate through the environment and reach the receive antennas via different paths, including line-of-sight, reflections from walls and furniture, and reflections from the human body. The presence and motion of a human subject alter the signal correlations on time, frequency or space,, resulting in changes to key signal properties such as Received Signal Strength (RSS), Angle of Arrival (AoA), and Time of Flight (ToF). By analyzing these signal changes, it is possible to infer meaningful information about human activities. Both handcrafted feature-based methods and data-driven deep learning models have been developed to map variations in Wi-Fi signals to human presence, location, motion, respiration, and more—thus enabling a wide range of Wi-Fi sensing capabilities.
@@
Figure 2: Wi-Fi sensing generalization is primarily hindered by three key factors, i.e., device heterogeneity, human body diversity, and environmental diversity.
Figure 3: Growth of research in Wi-Fi generalization: from a handful of studies between year of 2015 and 2018 to a surge of publications since 2019.
However, one critical challenge remains unresolved: generalization—the performance of Wi-Fi sensing often degrades when applied to new users, devices, or environments due to significant domain shifts. This performance gap between research prototypes and practical systems arises from the inherent variability in Wi-Fi sensing conditions, which introduces considerable domain shifts across deployment scenarios. As illustrated in Fig. 2, generalization is primarily hindered by three key factors, including device heterogeneity, human body diversity, and environmental diversity. First, device heterogeneity introduces substantial signal variation. Wi-Fi signals originate from a wide range of devices—including routers, laptops, and mobile phones—that differ in protocol versions, operating channels, bandwidth, antenna design, physical orientation, and chipset configurations. These variations affect the measured CSI and RSS, often rendering models trained on one device ineffective on another. Second, human body diversity introduces non-negligible discrepancies in signal interactions. Variations in height, weight, body composition, and even clothing materials can alter the propagation and absorption of Wi-Fi signals. As a result, two individuals performing the same action at the same location may produce significantly different signal patterns, posing challenges to consistent activity recognition. Third, environmental diversity in indoor scenes further exacerbates generalization difficulties. Differences in room layouts, wall materials, furniture placement, and device positions reshape the signal propagation paths and multipath profiles, leading to domain-specific variations in signal features such as RSS, AoA, and ToF. These changes severely limit the Wi-Fi sensing systems from generalizing beyond their training conditions.
@@
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=6em,align=left,fit=band,font=,, where level=2text width=8.2em,align=left,fit=band,font=,, where level=3text width=9em,font=,, where level=4text width=8em, font=,, [ Wi-Fi Human Sensing Generalization Categories, ver [ Experimental
Setup
II-A) , fill=orange!5 [ Distributed Antennas
II-A1), fill=orange!10,text width=9em [ WiNDR [63], WiCross [64], WiSDAR [65] , leaf,fill=orange!10, text width=42em ] ] [ Distributed Devices
II-A2),fill=orange!10,text width=9em [ ReWiS [66], Widar 3.0 [67], OneFi [68], WiTraj [69] , leaf,fill=orange!10, text width=42em ] ] [ Scaling Up Dataset
II-A3),fill=orange!10,,text width=9em [ XRF55 [70], Widar 3.0 [67], Zhang et al.[71], MM-Fi[72], CSI-Bench [73] , leaf,fill=orange!10, text width=42em ] ] ] [ Signal
Preprocessing
II-B), fill=violet!5 [ Signal Analysis
Methods (§II-B1), fill=violet!10,text width=9em [ WiHand [74], CLAR [75], Gait-Enhance[76], Zhuo et al.[77], Abuhoureyah et al. [78], leaf, fill=violet!15, text width=42em ] ] [ Advanced Handcrafted
Indicators (§II-B2), fill=violet!10, text width=9em [ WiHGR [79], PhaseAnti [80], DPSense [81], HandGest [82], Yu et.al. [83], AF-ACT [84], RoSeFi [85],
Wang et.al. [86], WiGesFree[87], LiKey [88], Bu et.al. [89], Zhang et.al. [90] , leaf, fill=violet!15, text width=42em ] ] [ Motion Indicators
II-B3), fill=violet!10,text width=9em [ Sharp [91], OneSense [92], Yin et al.[93], Wi-learner [94], Niu et al. [95], and LAGER [96], WiLife [97],
Mini-Batch Alignment [98, 99], WiTraj [69], Widar 3.0 [67]. Zhang et al. [100], WiTransformer [101],
Shi et al. [102], Bulugu [103]. GaitID [30], OneFi [68], XFall [104], WiGesture [105], Shi et al. [106],
Wang et al. [107], Widir [108], Wang et al. [109], WiNDR[63], WiDIGR [110], Wi-PIGR [111],
WiDetect [112], Zhu et al. [113], leaf, fill=violet!15, text width=42em ] ] [ Angle Indicators
II-B4), fill=violet!10,text width=9em [ Li et al. [114], EasyTrack [115], 3D-ID [116], leaf, fill=violet!15, text width=42em ] ] ] [ Feature
Learning
II-C),fill=yellow!5 [ Domain Alignment
II-C1),fill=yellow!10,text width=9em [ WiHARAN [117], DA-HAR [118], Zhang et al. [119], i-Sample [120], JADA [121], CrossGR [122], Wi-Adaptor [123],
EI [124], Khattak [125], WiCAR [126], GESFI [127], ADA [128], DATTA [129], WiSR [130], Yin et al.[93],
INDG-Fi [131], WiCross [64], Berlo et al. [132], DINN [133], freeGait [134], DAFI [135], Shi et al. [106], Fidora [136],
MMD [137], AdaPose [138], BSWCLoc [139], Kang et al. [140], WiAi-ID [141], Wi-CHARS [142], Mehryar [143],
AdapLoc [144], PIAS [145], CDD [146], WiCAU [147], WiSDA [148], FDAS [149], CDFi [150], LAGER [96],
Zhan & Wu [151], AirFi [152], WiLISensing [153], MetaGanFi [154], Zhang et al. [155] , Wi-Cro [156], WiTeacher [157],
CrossSense [158], Sugimoto et al. [159], CSI-MTGN [160], Wi-Fringe [161], XRF55 [70], Wi-Chat [162], WiFi2Radar [163] , leaf,fill=yellow!15, text width=42em ] ] [ Component Disentangle
II-C2),fill=yellow!10, text width=9em [ Elujide et al. [164], UH-Sense [165], Person-in-WiFi [45], Wi-Piga [166], CrossID [167], LT-WiOB [168], CATFSID [169],
TransferSense [170], UniFi [171], WiOpen [172], NCA [173], SFTSeg [174], DualConFi[175], CFLH [176], Xiao et al. [177],
3D-ID [116], MaP-SGAN [178], CrossFi [179], Wang et al. [180], WiGR [181], LESS [182], WiGesID [183] , leaf,fill=yellow!15, text width=42em ] ] [ Metric Learning for
Zero/Few-shot Learning
II-C3),fill=yellow!10, text width=9em [ LT-WiOB [168], CATFSID [169], TransferSense [170], UniFi [171], WiOpen [172], NCA [173], SFTSeg [174],
DualConFi[175], CFLH [176], Xiao et al. [177], 3D-ID [116], MaP-SGAN [178], CrossFi [179], Wang et al. [180],
WiGR [181], LESS [182], WiGesID [183], AFEE-MatNet [184], HAR-MN-EF [185], MatNet-eCSI [186], Zhou et al. [187],
WiLiMetaSensing [188], ReWiS [66], Ding et al. [189], WiONE [190], CAUTION [191], FewCS [192], Zhang et al. [193] , leaf,fill=yellow!15, text width=42em ] ] [ Meta Learning for
Few-shot Learning
II-C4),fill=yellow!10,text width=9em [ MAML [194], Huang et al. [195], TOSS [196], MetaLoc [197], MetaFormer [198], OneSense [92], Wi-learner [94],
RF-Net [199], WiGR [200], MetaGanFi [154], Zhang et al. [201] , leaf,fill=yellow!15, text width=42em ] ] [ Data Augmentation
and Data Synthesize
II-C5),fill=yellow!10,text width=9em [ CrossGR [122], Wang et al. [180], SIDA [202], WiARGAN [203], AFSL-HAR [204], LESS [182], Wi-Dist [205],
CsiGAN [206], Fidora [136], Fido [207], freeLoc [208], freeGait [134], Xiao et al. [177], Zhou et al. [209], CSI-Net [210],
BullyDetect [211], P3ID [212], OpenFi [213], Cutout [214], AirFi [152], RFBoost [215], WiSGP [216], WiAG [217],
Wi-Cro [156], SFTSeg [174], OneFi [68], OneSense [92], Wang et al. [218] , leaf,fill=yellow!15, text width=42em ] ] [ Pseudo Labeling
II-C6),fill=yellow!10,text width=9em [ CLAR [75], LAGER [96], Zhang et al. [219], DA-HAR [118], Zhang et al. [100] , leaf,fill=yellow!15, text width=42em ] ] ] [ Model
Deployment
II-D), fill=teal!3 [ Transfer Learning
II-D1), fill=teal!7,text width=9em [ XRF55 [70], BullyDedect [211], WiLISensing [153], DADA-AD [220], Brinke & Meratnia [221], WiTransfer [222],
FewSense [223], ResMon [224], RoGER [225], Hou et al. [226], DASECount [227] , leaf, fill=teal!10,, text width=42em ] ] [ Federated Learning
II-D2), fill=teal!7,text width=9em [ WiFederated [228], pFedBKD [229], CARING [230], Qi et al. [231], Co-WiSensing [232], Zhang et al. [119],
AdaWiFi [233], Zhang et al. [234] , leaf, fill=teal!10,, text width=42em ] ] [ Continual Learning
II-D3), fill=teal!7,text width=9em [ FewCS [192], WiDrive [235], M-WiFi [236], RISE [237], CCS [238], CAREC [239], CSI-ARIL [240] , leaf, fill=teal!10, text width=42em ] ] ] ]
@@
Figure 4: We systematically review and categorize nearly a decade of research on Wi-Fi sensing generalization, covering over 200 publications since 2015. Distinct from previous surveys, we organize the literature along a four-stage sensing pipeline—device deployment, signal preprocessing, feature learning, and model deployment.
@@
To enhance the generalization of Wi-Fi sensing systems, a diverse array of techniques has been proposed. As shown in Fig. 3, research in this field has expanded significantly, evolving from a few studies published between 2015 and 2018 to a surge of recent work focused on improving Wi-Fi generalization. These efforts can be broadly classified into four key stages of Wi-Fi sensing systems: experimental setup, signal preprocessing, feature learning, and model deployment.
@@
Figure 5: In the device deployment stage, Wi-Fi generalization can be enhanced by distributing antennas to mitigate the impact of user orientation, deploying devices more widely or optimizing their placement to increase coverage, and collecting more diverse datasets to support the training of more robust and generalizable models.
This survey systematically reviews and categorizes nearly a decade of research on Wi-Fi sensing generalization, covering over 200 publications since 2015, as shown in Fig. 4. Distinct from previous surveys [243, 244], we organize the literature along a four-stage sensing pipeline—experimental setup, signal preprocessing, feature learning, and model deployment—to provide a structured and comprehensive view of existing techniques. This stage-wise organization enables researchers and practitioners to quickly identify methods suited to their specific challenges. Furthermore, we curate and summarize a wide array of publicly available datasets, offering guidance for researchers seeking appropriate benchmarks to train or evaluate generalizable models. To support future research, we also explore emerging trends, including multimodal sensing, large pre-trained models, and data-efficient learning paradigms, that are poised to shape the next generation of Wi-Fi sensing technologies. We hope this survey will serve as a comprehensive and practical resource for both newcomers and experienced researchers seeking to understand, implement, or advance generalization techniques in Wi-Fi sensing.
@@
The remainder of this paper is organized as follows. Section II provides a stage-wise taxonomy of generalization techniques in Wi-Fi sensing, covering key methods and representative studies across device deployment, data preprocessing, feature learning, and model deployment, respectively. Section III presents a comprehensive overview of publicly available Wi-Fi sensing datasets, categorized by sensing tasks, including action recognition, pose estimation, crowd counting, etc. Section IV discusses promising future research directions, including multimodal sensing, large-scale model pretraining, and continual learning. Finally, Section V concludes the paper by summarizing key insights on the path forward for building more generalizable Wi-Fi sensing systems.
@@ II Taxonomy
To enhance the generalization of Wi-Fi sensing systems, researchers must address challenges across various stages, including device deployment, signal processing, feature learning, and model deployment. While some studies concentrate on a single stage, others tackle multiple stages simultaneously, addressing two or three phases at once. To provide a clear and systematic overview of existing research, we categorize the surveyed works based on the Wi-Fi sensing pipeline, as illustrated in Fig. 4.

### v4

@@ Abstract
Wi-Fi sensing has emerged as a powerful non-intrusive technology for recognizing human activities, monitoring vital signs, and enabling context-aware applications using commercial wireless devices. However, the performance of Wi-Fi sensing often degrades when applied to new users, devices, or environments due to significant domain shifts. To address this challenge, researchers have proposed a wide range of generalization techniques aimed at enhancing the robustness and adaptability of Wi-Fi sensing systems. In this survey, we provide a comprehensive and structured review of over 200 papers published since 2015, categorizing them according to the Wi-Fi sensing pipeline: experimental setup, signal preprocessing, feature learning, and model deployment. We analyze key techniques, including signal preprocessing, domain adaptation, meta-learning, metric learning, data augmentation, cross-modal alignment, federated learning, and continual learning. Furthermore, we summarize publicly available datasets across various tasks, such as activity recognition, user identification, indoor localization, and pose estimation, and provide insights into their domain diversity. We also discuss emerging trends and future directions, including large-scale pretraining, integration with multimodal foundation models, and continual deployment. To foster community collaboration, we introduce the Sensing Dataset Platform (SDP) (http://www.sdp8.org/) for sharing datasets and models. This survey aims to serve as a valuable reference and practical guide for researchers and practitioners dedicated to improving the generalizability of Wi-Fi sensing systems. Notably, while this paper focuses on Wi-Fi sensing, it is important to emphasize that the methodologies discussed in the feature learning and model deployment stages, e.g., domain alignment, metric learning, meta-learning, federated learning, and continual learning, are equally applicable to a broader range of wireless sensing generalization challenges, such as those encountered in millimeter radar-based human sensing. Given the rapid evolution of this field, we will continuously maintain and update relevant resources at https://github.com/aiotgroup/awesome-wireless-sensing-generalization.
@@
Wi-Fi sensing, Human sensing, Action recognition, Indoor localization, Domain adaptation, Domain generalization, Few-shot learning, Transfer learning, Meta-learning
@@ I Introduction
While Wi-Fi sensing has shown great promise across a variety of applications, its effectiveness fundamentally relies on how wireless signals interact with the surrounding environment and the human body. Fig. 1 illustrates how human presence and motion influence the propagation of Wi-Fi signals. As shown in the figure, Wi-Fi devices are typically equipped with multiple transmit and receive antennas. Transmit antennas emit radio signals that propagate through the environment and reach the receive antennas via different paths, including line-of-sight, reflections from walls and furniture, and reflections from the human body. The presence and motion of a human subject alter the signal correlations on time, frequency or space, resulting in changes to key signal properties such as Received Signal Strength (RSS), Angle of Arrival (AoA), and Time of Flight (ToF). By analyzing these signal changes, it is possible to infer meaningful information about human activities. Both handcrafted feature-based methods and data-driven deep learning models have been developed to map variations in Wi-Fi signals to human presence, location, motion, respiration, and more, thus enabling a wide range of Wi-Fi sensing capabilities.
@@
Figure 2: Wi-Fi sensing generalizability is primarily hindered by three key factors, i.e., device heterogeneity, human body diversity, and environmental diversity.
Figure 3: Growth of research in Wi-Fi sensing generalizability: from a handful of studies between year of 2015 and 2018 to a surge of publications since 2019. We employed a linear regression to fit the growth trend. In the figure, the $R^{2}$ score indicates the correlation between the estimated results and the ground truth, where a value closer to 1 denotes a higher degree of model fitting. (We selected one representative paper per year based on the highest citation count, ensuring each choice employs a technical approach distinct from those featured in previous years.)
However, one critical challenge remains unresolved: Wi-Fi sensing generalization—the overall capability of Wi-Fi sensing systems to perform effective inference in unseen domains, such as new users, devices, or environments. To ensure clarity, the related term Wi-Fi sensing generalizability in this survey denotes the inherent property of a Wi-Fi sensing system design that enables such adaptability.
This performance gap between research prototypes and practical systems arises from the inherent variability in Wi-Fi sensing conditions, which introduces considerable domain shifts across deployment scenarios. As illustrated in Fig. 2, generalization is primarily hindered by three key factors, including device heterogeneity, human body diversity, and environmental diversity. First, device heterogeneity introduces substantial signal variation. Wi-Fi signals originate from a wide range of devices, including routers, laptops, and mobile phones, that differ in protocol versions, operating channels, bandwidth, antenna design, physical orientation, and chipset configurations. These variations affect the measured CSI and RSS, often rendering models trained on one device ineffective on another. Second, human body diversity introduces non-negligible discrepancies in signal interactions. Variations in height, weight, body composition, and even clothing materials can alter the propagation and absorption of Wi-Fi signals. As a result, two individuals performing the same action at the same location may produce significantly different signal patterns, posing challenges to consistent activity recognition. Third, environmental diversity in indoor scenes further exacerbates generalization difficulties. Differences in room layouts, wall materials, furniture placement, and device positions reshape the signal propagation paths and multipath profiles, leading to domain-specific variations in signal features such as RSS, AoA, and ToF. These changes severely limit the Wi-Fi sensing systems from generalizing beyond their training conditions.
@@
Figure 4: We systematically review and categorize nearly a decade of research on Wi-Fi sensing generalization, covering over 200 publications since 2015. Distinct from previous surveys, we organize the literature along a four-stage sensing pipeline—experimental setup, signal preprocessing, feature learning, and model deployment.
@@
To enhance the generalizability of Wi-Fi sensing systems, a diverse array of techniques has been proposed. As shown in Fig. 3, research in this field has expanded significantly, evolving from a few studies published between 2015 and 2018 to a surge of recent work focused on improving Wi-Fi generalization. These efforts can be broadly classified into four key stages of Wi-Fi sensing systems: experimental setup, signal preprocessing, feature learning, and model deployment.
@@
TABLE I: Common Abbreviations in the Article for Quick Retrieval.
Abbreviation Full Name
ADoA Angular Difference of Arrival
AoA Angle of Arrival
AoD Angle of Departure
AP Access Point
BFI Beamforming Feedback Information
BFM Beamforming Feedback Matrix
BVP Body-coordinate Velocity Profile
CNN Convolutional Neural Network
CSI Channel State Information
CWT Continuous Wavelet Transform
DFS Doppler Frequency Shift
DPV Dynamic Phase Vector
GAN Generative Adversarial Networks
ICA Independent Component Analysis
ISAC Integrated Sensing and Communications
MAML Model-Agnostic Meta-Learning
MMD Maximum Mean Discrepancy
NLOS Non-Line-of-Sight
PCA Principal Component Analysis
RKHS Reproducing Kernel Hilbert Space
RNN Recurrent Neural Network
RSS Received Signal Strength
RSSI Received Signal Strength Indicator
STFT Short-Time Fourier Transform
ToF Time of Flight
VAE Variational Autoencoder
@@
Figure 5: In the experimental setup stage, Wi-Fi generalization can be enhanced by distributing antennas to mitigate the impact of user orientation, deploying devices more widely or optimizing their placement to increase coverage, and collecting more diverse datasets to support the training of more robust and generalizable models.
TABLE II: Comparison of Experimental Setup Strategies for Enhanced Generalization.
Category Pros (Advantages) Cons (Limitations) Generalization Mechanism Best-fit Scenarios
Distributed Antennas ([165, 164, 202]) Achieves direction-agnostic sensing with single device pairs; minimal additional hardware cost. Requires manual antenna placement; limited coverage in large-scale areas. Introduces angular diversity by capturing signals from multiple viewpoints, reducing sensitivity to human orientation and body aspect angle. Gesture and activity recognition tasks where user orientation varies but deployment scale is limited.
Distributed Devices ([9, 313, 244, 235]) Provides high spatial resolution; reduces blind spots and mitigates multipath occlusions. High deployment cost; complex synchronization and calibration across devices. Exploits spatial diversity by observing heterogeneous signal responses across locations, improving robustness to environmental variations. Large indoor spaces or complex layouts requiring robust cross-location generalization.
Scaling Up Dataset ([208, 313, 304, 254, 323]) Enables models to learn implicit invariances; improves robustness and supports few-shot transfer. Labor-intensive data collection; increased storage and training computation cost. Enhances population-level generalization by exposing models to diverse human body types, motion patterns, and interaction styles. Data-driven systems targeting user-independent sensing across diverse populations and activities.
In the literature, there have been two recent surveys focusing on Wi-Fi sensing generalization. Chen et al. [24] adopt a cross-domain perspective and review over 60 representative studies, emphasizing learning-based techniques such as domain-invariant feature extraction, transfer learning, and few-shot learning. Wang et al. [229] focus more narrowly on the few-shot learning paradigm, providing an in-depth discussion of metric learning, transfer learning, and meta-learning approaches for data-scarce scenarios.
While these surveys provide valuable insights into specific aspects of Wi-Fi sensing generalization, they primarily concentrate on learning paradigms and do not offer a system-level view of how generalization challenges arise and are addressed across the entire sensing workflow. In contrast, our survey reviews and synthesizes over 200 publications since 2015 and provides several new insights beyond prior work. First, we organize existing techniques along a four-stage Wi-Fi sensing pipeline, i.e., experimental setup, signal preprocessing, feature learning, and model deployment, revealing how generalization challenges and solutions manifest differently at each stage. Second, to provide actionable guidance, we include “Key Takeaways and Lessons Learned” for each stage. These sections offer a comparative analysis of different techniques’ pros, cons, generalization mechanisms, and best-fit scenarios. Third, by jointly analyzing tools, datasets, learning paradigms, and deployment considerations, our survey moves beyond current techniques to propose a structured research roadmap with short-term and long-term goals for emerging fields like Wi-Fi foundation models and large multimodal models. By providing these multi-dimensional insights—from fundamental principles to practical deployment decisions and future strategic planning—our work offers a more holistic and instructive framework that complements and significantly extends existing literature.
@@
The remainder of this paper is organized as follows. Section II provides a stage-wise taxonomy of generalization techniques in Wi-Fi sensing, covering key methods and representative studies across experimental setup, data preprocessing, feature learning, and model deployment, respectively. Section III presents a comprehensive overview of publicly available Wi-Fi sensing datasets, categorized by sensing tasks, including action recognition, pose estimation, crowd counting, gait recognition, etc. Section IV discusses promising future research directions, including multimodal sensing, large-scale model pretraining, and continual learning. Finally, Section V concludes the paper by summarizing key insights on the path forward for building more generalizable Wi-Fi sensing systems. An abbreviation table is provided in Table I, showcasing the abbreviations of core Wi-Fi signals or methods for quick reference.
@@ II Taxonomy
To enhance the generalization of Wi-Fi sensing systems, researchers must address challenges across various stages, including device deployment, signal processing, feature learning, and model deployment. While some studies concentrate on a single stage, others tackle multiple stages simultaneously, addressing two or three stages at once. To provide a clear and systematic overview of existing research, we categorize the surveyed works based on the Wi-Fi sensing pipeline, as illustrated in Fig. 4.